In an intelligence test conducted on 1000 students, the average scores was 42 and standard deviation 24. 

Find the number of successful students if the minimum pass marks is 40.


1
Expert's answer
2021-05-18T14:22:10-0400

Let "X=" the student's mark: "X\sim N(\mu, \sigma^2)"

Given "\mu=42, \sigma=24"


"P(X\geq40)=1-P(X<40)"

"=1-P(Z<\dfrac{40-42}{24})\approx1-P(Z<-0.083333)"

"\approx1-0.4668=0.5332"




"0.5332\cdot1000=533"

The number of successful students is 533.